A hose can fill a children's pool in 3 hours. If the drain is open, the pool will drain in 5 hours. If the hose is turned on AND the drain is left open, how long will it take to fill the children's pool?

Answer by stanbon(57290)

A hose can fill a children's pool in 3 hours. If the drain is open, the pool will drain in 5 hours. If the hose is turned on AND the drain is left open, how long will it take to fill the children's pool?
--------------------------
fill rate = 1/3 job/hr
drain rate = 1/5 job/hr
together rate = 1/x job/hr
------------------------------
Equation:
fill - drain = 1 job
1/3 - 1/5 = 1/x
---
5x - 3x = 15
2x = 15
x = 7.5 hrs. (time to fill the pool)
=======================================
Cheers,
Stan H.
============